Scenario: Income-Expenditure Equilibrium GDP is $8000, autonomous consumption is $500, and planned investment spending is $20Why is the answer -$100?
Answer #1

an 4 where A bar is equal to autonomous consumption and planned investment spending.

Now change in Y is 3000- 8000 = -5000

1/1-c = 1/1-0.8 = 1/ 0.2 = 5

Change in Y/ change in A bar = 5

-5000/Change in A bar = 5

Change in A bar = -5000/5

= -1000

It is -100 because if it is -100, new autonomous spending ( A bar) will be 500 + 100

Change in Autonomous spending = 600-700

= -100